# Rethinking the Atmospheric Scattering-driven Attention via Channel and Gamma Correction Priors for Low-Light Image Enhancement

### v3

@@ III-B Relinking the Gamma Correction to Global-Local Processing
Building on insights from [10], we understand that the adaptive gamma value serves as an environmental factor representing the overall illumination conditions, which can vary across different scenes. Consequently, this attention module ensures that the network focuses on regions within the broader image where gamma correction yields the most significant enhancement. This approach is called “Plug-in Attention with Gamma Correction,” and the CP block with IAAF is named “CPGA block.”
III-C Loss Functions
@@
In supervised learning, we use four loss functions to guide our approach: L1 loss, perceptual loss, HDR L1 loss, and SSIM loss. We utilize HDR L1 loss, perceptual loss, and total variation loss for unsupervised learning, with the target set as the histogram-equalized low-light image.
@@
The L1 loss function, a commonly use loss function that performs better in image enhancement and restoration, is defined as:
@@
HDR L1 loss, as introduced in [23] is computed in the tone-mapped domain since HDR (High Dynamic Range) images are typically viewed after tone-mapping. To achieve this, they apply the widely used µ-law function to calculate the loss:
@@
where $\mu$ is set to 5000, $T$ is the tone-mapped HDR image, and $x$ is the input image. Then, the $\mu$-law function is applied to the L1 loss as follows:
@@
$L_{\text{SSIM}}=1-\text{SSIM}(\hat{Y},Y^{\text{GT}})$ (16)
For unsupervised learning of our approach, the only optimal for brightening is the histogram-equalized low-light image, which is a good aim for the contrast but not for the details. Thus, we utilized total variation loss [24, 25] for denoising and improving the smoothness:
@@
III-D Revealing the Explainable Mechanisms Behind Degradation
Our method follows a rule-based learning strategy, leveraging the ATSM and the Retinex theory to perform image processing based on physical models. The ATSM simulates the scattering and absorption of light in uneven media, while the Retinex theory, inspired by retinal imaging, models the human visual system. Although these methods provide solid theoretical support from both physical and physiological perspectives, challenges remain in the interpretability of neural networks. While applying the models to deep learning, it is difficult to explain its decision-making mechanisms due to the highly nonlinear nature of deep learning models. Similarly, these are hard to define and are barely explained through experience or a simple understanding of the equations from the generated features in the neural network. To improve the interpretability of our approach, we utilize the traditional image degradation model, which offers a more explicit theoretical framework for the deep learning process. This effectively enhances the transparency of the entire neural network, opening the “black box” of image enhancement.
Initially, the fundamental image degradation model in matrix form [1] can be defined as:
[…]
Figure 6: Visualization of extracted features corresponding to Eq. (3) and (11) [8]. (a) Low-light image $L$; (b) a portion of $L$; (c) $\tilde{t}$ (contribution proportion from $L$); (d) An estimated image of $\tilde{A}$; (e) Reconstructed low-light image $R$; (f) Gamma-corrected image $R^{\gamma}$; (g) Enhanced image $\hat{R}$.
We attempt to further verify our hypotheses for the local branch through least-squares optimization [26]. First, we make the following simple assumptions: we assume that $\hat{A}$ and $L$ can be expressed as the noise-free reconstructed image $R^{\text{GT}}$ with additional composite noises that consist of three-channel mixing factors, lighting changes, and inherent thermal noises, which are particularly prominent in low-light environments, as depicted below:
$\tilde{A}=R^{\text{GT}}+N_{A}$ (20)
$L=R^{\text{GT}}+N_{L}$ (21)
where $L$ is the low-light image, $N_{A}$ and $N_{L}$ denote the noise components associated with $\tilde{A}$ and $L$, respectively. While estimating $N_{A}$, we assume $N_{L}$ is fixed once the low-light image $L$ is given. Then, we can define the following residual error $r(\tilde{t},N_{A})$ by combining Eq. (3), Eq. (20), and Eq. (21):
$r(\tilde{t},N_{A})=R-R^{\text{GT}}=(1-\tilde{t})N_{A}+\tilde{t}N_{L}$ (22)
We further define a cost function $C(\tilde{t},N_{A})$ for the least-square optimization process:
$C(\tilde{t},N_{A})=\frac{1}{2}\left[r(\tilde{t},N_{A})\right]^{2}$ (23)
$\frac{\partial C}{\partial N_{A}}=r(\tilde{t},N_{A})\cdot\left(-\tilde{t}\right)=0$ (25)
Case I: $r(\tilde{t},N_{A})\neq 0$: We have
$(N_{A}-N_{L})=0\quad\text{and}\quad\tilde{t}=0$ (26)
which leads to the solution $\tilde{t}=1$ and $N_{L}=N_{A}$, resulting in:
$r_{\text{min}}(\tilde{t},N_{A})=N_{A}$ (27)
However, $\tilde{t}=0$ results in $R=L$ from Eq. (3), meaning that no restoration effect is involved at all (identity mapping from the input $L$ to the output $R$).
Case II: $r(\tilde{t},N_{A})=0$: We have
$R-R^{\text{GT}}=0\implies R=R^{\text{GT}}$ (28)
and
Therefore, the optimal solution exists when Eq. (29) holds.
@@ IV-B Implementation Details
Our experiments are conducted and evaluated with an NVIDIA GeForce RTX 3090 GPU. For the preprocessing of training, we cropped the image into $256\times 256$ pixels with random translation. We set 600 epochs for training. An initial learning rate of $10^{-3}$is set for training with the Adam optimization scheme, and the learning rate changing cycle is 67 epochs by the Cosine Annealing scheduler. To address computational demand, we implemented simplified block pruning named “Block Design Simplification (BDSF)” in our lightweight version, directly removing blocks in the local branch for inference. Furthermore, we have also applied our approach as unsupervised learning using histogram equalization, demonstrating its flexibility. When using unsupervised learning, we trained on LOLv2 Real for 50 epochs, which consists of more realistic data and does not require any normal exposure images for training.
@@ IV-C Evaluation Results
As shown in Tables I and II, our approach reaches a high standard compared to others, which ranked third on paired datasets and ranked first on five sets of unpaired data, while the model remains lightweight design with low numbers of parameters and FLOPs. The visualized comparisons are shown in Figs. 7 and 8. Also, we figured out that the approaches with CPGA architecture present better quality on unpaired data, which means that our theoretical equations assumptions for improvement are closer and related to nature, making the image more realistic. Furthermore, our method successfully improves the performance via the same architecture by 5% SSIM compared to the CPGA-Net while maintaining the lightweight design.
For the comparison of unsupervised approaches, as shown in Table II we ranked first compared to other unsupervised approaches with better contrast. This demonstrates the robustness of our theoretical-based network architecture when using simple supervision of histogram-equalized images. However, there are more noticeable defects and distortions due to the lack of strong supervision of the details, as shown in Fig. 9. This will be a focus for our future work.
@@
Original Image and Method Types MEF LIME NPE VV DICM Avg
@@
Figure 7: Visual comparison on paired datasets [5].
Figure 8: Visual comparison on unpaired datasets [11, 27, 30].
@@ IV-D High-Level Vision Task
In this section, we address the challenge of objection detection in low-light environments by utilizing a joint training approach of YOLOv9s [33] with the SOTA approaches of LLIE on the ExDark dataset [31], as illustrated in Table III. Our approach improves the mean Average Precision (mAP) by 0.075 compared to baseline. All the LLIE methods listed here can improve object detection performance, among which the Zero-DCE [13] and our proposed method are the best. The results show that our proposed method can improve not just human perception but machine perception as well.
@@ V Ablation Study
In this section, we analyze the effectiveness of each systematic module and training technique, including the systematic design, the number of Channel-Prior blocks, and loss functions.
@@ V-A Systematic Design and Integration
As shown in Table IV, our method effectively fuses the gamma correction from the global branch to the local branch, resulting in improved overall performance and demonstrating the strength of our approach. By grounding the attention mechanism in gamma correction, we ensure that the enhancement process remains aligned with the non-linearities inherent in both the imaging process and human perception.
On the other hand, we also figured out that the global branch alone performs surprisingly well. In Section III-D, we elaborate on the system underlying the equations and highlight the relationship between $R$ and $R^{\text{GT}}$, and the results suggest that this correlation is stronger than expected. For instance, Table IV(c) and Table IV(e) show only a minor difference in PSNR (0.5 dB), and there is no performance difference between Table IV(e) and Table IV(f) on LOLv1. However, despite the significant contribution of the global branch, Table IV(c) still shows a noticeable gap in achieving optimal performance without incorporating local processing. This underscores the necessity of the local branch to bridge the gap between local and global processing, ensuring the system achieves its highest potential in quality and consistency. Since our pruning approach is based on the concept of block pruning but differs from existing methods [34, 35, 36, 37] and retains its simplicity even without requiring additional training, we named the proposed technique ”Block Design Simplification (BDSF).”
@@
V-B The Numbers of Channel-Prior Blocks
We explore how varying the number of CP blocks affects the model’s capacity to enhance image quality. The results, summarized in Table V, show that increasing the number of CP blocks leads to an improvement from 0 to 2 blocks but show no significant changes from 2 to 4 blocks. However, both the number of parameters and computational cost (FLOPs) increase with more CP blocks, introducing greater computational demands. Therefore, the optimal number of CP blocks should balance performance gains with resource efficiency. For our final approach, we selected 2 CP blocks to achieve a lightweight and efficient design.
TABLE V: Ablation study of the number of CP blocks. $N=2$ is the default setting of our approach.
@@ V-C Loss Functions
This section examines the impact of various loss function combinations on the model’s performance. We tested L1 loss, Perceptual loss, HDR L1 loss, and SSIM loss, with the results summarized in Table VI. Using the default settings as in CPGA-Net, the combination of L1 and Perceptual losses performs well, yielding a PSNR improvement of 2.82 dB and an SSIM increase of 0.024. The HDR L1 loss significantly enhances all three metrics, with PSNR increasing by 2.86 dB, SSIM by 0.041, and LPIPS decreasing by 0.085. While SSIM loss improves its specific metric with an SSIM boost of 0.031, it is less effective in enhancing PSNR. Ultimately, combining all these losses results in the best overall performance for supervision, which improves PSNR by 4.04 dB, SSIM by 0.083, and LPIPS by 0.118.
TABLE VI: Ablation study of loss functions.

### v4

@@ III-B2 Plug-in Attention with Gamma Correction
Consequently, this attention module ensures that the network focuses on regions within the broader image where gamma correction yields the most significant enhancement. This approach is called “Plug-in Attention with Gamma Correction,” and the CP block with IAAF is named “CPGA block.”
III-C Exploiting Implicit Regularization of CPGA-Net
While our prior work, CPGA-Net [8], demonstrated strong performance by adhering to a theoretical image restoration model, a deeper analysis of its internal mechanisms reveals a critical limitation. As illustrated in Fig. 3, the feature maps generated by the local branch are nearly identical to the input, indicating that the network learns to approximate a near-identity mapping in this path. This situation creates redundancy and suggests the system fails to provide localized feature corrections. The main reason came from the design of the fusion strategy.
(a) PSNR
(b) SSIM
Figure 5: The learning curve comparisons of ablation results for CPGA-Net [8]. Blue represents the local branch only, red represents the global branch only, and green represents the complete CPGA-Net.
IAAF is an effective fusion strategy inspired by set theory, designed to blend two feature sets by compensating for their union with a learned intersection. Our primary objective was to employ IAAF to integrate the base reflectance $R$ and its gamma-corrected counterpart $R^{\gamma}$. The global branch manages brightness control, and the local branch concentrates on detail reconstruction. Our analysis in Fig. 5 revealed a significant issue with asynchronous convergence between the two branches. The global branch handles brightness adjustment through factorized gamma correction, which enhances contrast but often fails to capture fine details. In contrast, the local branch emphasizes restoring structural elements and delivers superior performance, although it presents greater training challenges.
While merging branches can yield a balanced quality, this discrepancy creates a training imbalance. Consequently, the final output of the local branch tends to be a suboptimal compromise, with global processing dominating the restoration process, leading to a critical challenge, as the dominance of the global branch suppresses the local branch, stabilizing the restoration but ultimately impeding overall performance. This suppression results in a near-identity mapping of the local branch, even while achieving improved outcomes, as shown in Fig. 6.
Figure 6: Visual analysis of the intermediate feature maps and fusion strategies in CPGA-Net+ and CPGA-Net++.
This analysis uncovers a fascinating and counterintuitive dynamic: an adversarial yet ultimately symbiotic relationship between the global and local branches. While converging to a near-identity mapping, the local branch introduces a subtle disturbance to the feature space. Crucially, instead of being a flaw, this disturbance acts as a powerful form of implicit regularization, supporting the global branch by forcing it to learn a more robust and generalized restoration mapping. Consequently, the entire restoration workload becomes centralized within the global branch, which, strengthened by this adversarial process, achieves a more stable and effective final solution. This regularization dynamic is crucial during training, but once the model is trained, the local branch’s function reverts to a simple, computationally redundant identity mapping during inference.
Therefore, we proposed a structural pruning technique based on the concept of block pruning by directly removing the local branch while inferencing because of this characteristic, but it differs from existing methods [23, 24, 25, 26] and maintains its simplicity without requiring any additional training. We named this technique “Block Design Simplification (BDSF).”
III-D Unleashing the Power of Intersection-Aware Adaptive Fusion
In contrast to the pruning, we further explore the solution to address the mismatch of data flow to utilize the overall capacity hidden in the local branch. We introduce IAAF+, a novel module designed to control the fusion process intelligently. We actively apply gamma correction to the input $L$ to address training imbalances, allowing for independent gradient flow. Then, the module operates through two core mechanisms: Weighted Feature Selection (WFS), which models the intersection of features, and Intersection Representation Scoring (IRS), which supplies adaptive fusion weights.
The Weighted Features Selection (WFS) module acts as a dynamic scoring function. It analyzes features from both branches to determine their optimal blending ratio, outputting a single learnable scalar parameter $\alpha$. This value is then used to compute a baseline feature union $^{\hat{R}}_{\cup}$, as a weighted average of the two branches:
$\hat{R}_{\cup}=\alpha R+(1-\alpha)L^{\gamma}.$ (9)
By making the fusion weight $\alpha$ learnable, the WFS module allows the network to balance the contribution of each branch throughout the training process. This creates an adaptive fusion strategy that harmonizes the two branches instead of letting them compete.
Concurrently, the Intersection Representation Scoring (IRS) module explicitly models the shared, redundant information with more accurate and detailed mapping in the intersection of IAAF:
$\hat{R}_{\cap}=\cap(S_{R}\odot R,S_{L^{\gamma}}\odot L^{\gamma}),$ (10)
where $S_{R}$ and $S_{L^{\gamma}}$ are the spatial attention maps produced via $R$ and $L$, respectively. Each feature map is modulated via element-wise multiplication $\odot$ with the corresponding spatial attention maps, allowing the intersection to emphasize reliable regions while suppressing noise, resulting a cleaner and more informative shared representation of fusion. This allows the module to perform a selective and weighted intersection calculation, focusing only on each branch’s most reliable feature regions to define the shared information. This prevents noisy or less relevant features from corrupting the intersection estimate, leading to a more precise and meaningful compensation in the final fusion step.
Overall, the final output, $\hat{R}(x)$, is produced by first creating a weighted average of the two branches controlled by WFS, and then subtracting the intersection calculated by IFS. This entire process is captured in the following formulation:
$\hat{R}_{\text{IAAF+}}=\hat{R}_{\cup}-\hat{R}_{\cap}$, (11)
which separates the task of global balancing (via WFS’s $\alpha$) from specific feature compensation (via IRS’s intersection map), promotes a synergistic relationship between the branches for a more robust and detailed final image. The diagram of IAAF+ is shown in Fig. 7.
Our module has a two-phase architecture that creates a deliberate separation of concerns. The intermediate stages focus on self-enhancement, where the network iteratively refines the global feature representation $R$ by learning to adjust its illumination and contrast against its gamma-corrected version. The final stage then performs the critical heterogeneous fusion, injecting the fine-grained details from the separate local branch $L^{\gamma}$ only after the global feature $R$ has been fully stabilized and optimized. The overall analysis compared to baseline is shown in Fig. 6.
Figure 7: A block diagram of the IAAF+ module. $\alpha$ is the learnable scalar parameter to identify the importance of input features toward fusion by the WFS module. In contrast, the IRS module uses the spatial attention maps $S$ to evaluate the effectiveness of input signals for intersection representation.
III-E Loss Functions
In this section, we present the training methodology of our approach, which incorporates both supervised and unsupervised learning strategies. The unsupervised scheme is designed to induce pseudo-supervised behavior, thereby extending the applicability of our method to real-world scenarios without ground-truth supervision.
III-E1 Supervised learning
@@
In supervised learning, we use four loss functions to guide our approach: L1 loss, perceptual loss, SSIM loss, and HDR L1 loss.
@@
The L1 loss function, a commonly used loss function that performs better in image enhancement and restoration, is defined as:
@@
$L_{\text{SSIM}}=1-\text{SSIM}(\hat{Y},Y^{\text{GT}})$ (14)
To better handle HDR image content, we further adopt the HDR L1 loss. As introduced in [28], it is computed in the tone-mapped domain since HDR (High Dynamic Range) images are typically viewed after tone-mapping. To achieve this, they apply the widely used µ-law function to calculate the loss:
@@
where $\mu$ is set to 5000, $T$ is the tone-mapped HDR image, and $x$ is the input image. Then, $\mu$-law function, we utilized the L1 loss as follows:
@@
III-E2 Unsupervised training strategy
While our supervised learning approach yields effective image enhancement, a fully supervised method becomes infeasible when ground truth normal-light images are unavailable. To address this limitation, we propose an unsupervised training strategy aligned with our overall network framework, which effectively balances contrast, perceptual quality, and smoothness, offering a robust alternative when ground-truth supervision is unavailable. For our unsupervised learning process, we utilize HDR L1 loss, perceptual loss, and total variation loss, with the target being the histogram-equalized low-light image.
First, we use histogram-equalized images with HDR L1 and perceptual loss to adjust tone mapping and content. This guides the network toward producing outputs with appropriate tone mapping. Next, inspired by findings in [8], we supervise mid-feature maps using a blurred low-light input with HDR L1 loss. Thus, we utilized total variation loss [29, 30] for denoising and improving the smoothness:
@@ III-E2 Unsupervised training strategy
This unsupervised learning strategy enables our model to generalize to real-world low-light conditions by pseudo-supervised behavior without access to reference images.
@@ IV-B Implementation Details
Our experiments are conducted and evaluated with an NVIDIA GeForce RTX 3090 GPU. For the preprocessing of training, we cropped the image into $256\times 256$ pixels with random translation. We set 600 epochs for training. An initial learning rate of $10^{-3}$ is set for training with the Adam optimization scheme, and the learning rate changing cycle is 67 epochs by the Cosine Annealing scheduler. To manage computational demands, we introduced a method called “Block Design Simplification (BDSF)” in our lightweight version, which directly removes blocks in the local branch during inference. For CPGA-Net++, we utilized IAAF+ to replace IAAF [8] and substituted ResBlock [22] with ConvNeXT block [36]. Due to the ConvNeXT structure, the training of CPGA-Net++ was changed to AdamW with a cosine annealing scheduler until completion. The training of CPGA-Net++ also includes an intermediate supervision of 5 epochs to guide local processing toward the ground truth, ensuring effective learning for the local branch. We have also applied our approach as unsupervised learning using histogram equalization, demonstrating its flexibility. When using unsupervised learning, we trained on LOLv2 Real for 50 epochs, which consists of more realistic data and does not require any normal exposure images for training.
@@ IV-C Evaluation Results
Tables I and II show that, our approach achieves a higher standard than other methods, ranking second and third on paired datasets and first on five sets of unpaired data. Furthermore, our BDSF model maintains an ultra-lightweight design with a low number of parameters and FLOPs, resulting in only a minimal decrease in performance. Meanwhile, CPGA-Net++ is significantly enhanced by leveraging the potential of the local branch, thereby achieving strong quality and competitive performance. Visual comparisons are presented in Figs. 8 and 9. Also, we figured out that the approaches with CPGA architecture present better quality on unpaired data, which means that our theoretical equations’ assumptions for improvement are closer and related to nature, making the image more realistic. Furthermore, our approach improves performance by 0.05 SSIM over CPGA-Net using the same architecture, while maintaining a lightweight design. Notably, CPGA-based methods consistently yield better quality on unpaired data, suggesting that our equation-driven design aligns more closely with natural image properties, resulting in more realistic enhancements.
For the comparison of unsupervised approaches, as shown in Table II, we ranked first compared to other unsupervised approaches with better contrast. This demonstrates the robustness of our theoretically-based network architecture when using simple supervision of histogram-equalized images. However, there are more noticeable defects and distortions due to the lack of strong supervision of the details, as shown in Fig. 10. This will be a focus for our future work.
@@
Original Image and Method Types MEF LIME NPE $V$ DICM Avg
@@
CPGA-Net++ S 3.3825 2.8646 3.0434 1.9302 2.5241 2.7490
Figure 8: Visual comparison on paired datasets [5]. Zoom in for a better view.
Figure 9: Visual comparison on unpaired datasets [11, 31, 34]. Zoom in for a better view.
@@ IV-D High-Level Vision Task
In this section, we address the challenge of objection detection in low-light environments by utilizing a joint training approach of YOLOv9s [39] with the SOTA approaches of LLIE on the ExDark dataset [35], as illustrated in Table III. Our approach improves the mean Average Precision (mAP) by 0.075 compared to the baseline. All the LLIE methods listed here can improve object detection performance, among which the Zero-DCE [14] and our proposed method are the best. The results show that our proposed method can improve not just human perception but machine perception as well.
@@ V Ablation Study
In this section, we analyze the effectiveness of each systematic module and training technique, including the systematic design and integration, the number of Channel-Prior blocks, and loss functions.
@@ V-A Systematic Design and Integration
As shown in Table IV, our method effectively fuses the gamma correction from the global branch to the local branch, resulting in improved overall performance and demonstrating the strength of our approach. By grounding the attention mechanism in gamma correction, we ensure that the enhancement process remains aligned with the non-linearities inherent in the imaging process and human perception.
On the other hand, we received the surprising results for our lightweight (BDSF) and stronger pipeline (CPGA-Net++) designs, as shown in Tables IV and V. The BDSF strongly utilizes our model’s characteristics based on the fact that global processing dominates and suppresses the local processing, using an antagonism to maximize the power of global processing, leading to an ultimate and simple pruning to reduce the computational cost of local processing. In contrast, CPGA-Net++ employed IAAF+ to encourage both branches to operate effectively, resulting in improved performance. However, this approach diminishes the auxiliary function of local processing, making it impossible to achieve performance levels comparable to BDSF for the efficient version via pruning.
@@
TABLE V: Ablation study of structural functions for CPGA-Net++. The type indicates the fusion input source: Homo refers to homogeneous fusion using $R^{\gamma}$, while Hetero refers to heterogeneous fusion using $L^{\gamma}$. For the efficient version comparison, (b) is CPGA-Net+ (BDSF), and (i) is the finetune of CPGA-Net++ with the same structural pruning as BDSF, as CPGA-Net++ is not possible to directly prune while the characteristic of near-identity mapping is no longer maintained; therefore, we added a * beside the BDSF.
Block IAAF Efficient IQA Efficiency Resblock ConvNeXT Type WFS IRS PSNR↑ SSIM↑ LPIPS↓ # of P. (M) FLOPs (G) (a) Homo 22.53 0.812 0.205 0.060 9.356 (b) Homo BDSF 22.53 0.812 0.205 0.020 2.141 (c) Hetero 22.42 0.810 0.206 0.060 9.356 (d) Hetero 21.94 0.812 0.194 0.060 9.230 (e) Hetero 21.87 0.802 0.211 0.089 22.793 (f) Hetero 22.89 0.814 0.182 0.089 22.793 (g) Homo 21.34 0.799 0.231 0.089 22.793 (h) Hetero 22.24 0.835 0.136 0.062 13.302 (i) Hetero BDSF* 21.28 0.800 0.220 0.023 3.696
V-B The Number of Channel-Prior Blocks
We explore how varying the number of CP blocks affects the model’s capacity to enhance image quality. The results, summarized in Table VI, show that increasing the number of CP blocks leads to an improvement from 0 to 2 blocks but show no significant changes from 2 to 4 blocks. However, the number of parameters and computational cost (FLOPs) increase with more CP blocks, introducing greater computational demands. Therefore, the optimal number of CP blocks should balance performance gains with resource efficiency. We selected 2 CP blocks for our final approach to achieve a lightweight and efficient design.
TABLE VI: Ablation study of the number of CP blocks. $N=2$ is the default setting of our approach.
@@ V-C Loss Functions
This section examines the impact of various loss function combinations on the model’s performance. We tested L1 loss, Perceptual loss, HDR L1 loss, and SSIM loss, with the results summarized in Table VII. Using the default settings as in CPGA-Net, the combination of L1 and Perceptual losses performs well, yielding a PSNR improvement of 2.82 dB and an SSIM increase of 0.024. The HDR L1 loss significantly enhances all three metrics, with PSNR increasing by 2.86 dB, SSIM by 0.041, and LPIPS decreasing by 0.085. While SSIM loss improves its specific metric with an SSIM boost of 0.031, it is less effective in enhancing PSNR. Ultimately, combining all these losses results in the best overall performance for supervision, which improves PSNR by 4.04 dB, SSIM by 0.083, and LPIPS by 0.118.
TABLE VII: Ablation study of loss functions.